OpenStudy (anonymous):

Solve the following equation: (5x-1)/2+(-3x-1)/4=9/2 Can someone please show me how to work this out?

OpenStudy (precal):

\[\frac{5x-1}{2}+\frac{-3x-1}{4}=\frac{9}{2}\] is this your equation?